# The Three Regimes of Offline-to-Online Reinforcement Learning

### v1

@@ 4.1 Stability relative to the offline dataset $\mathcal{D}$
A common strategy for incorporating offline data during fine-tuning is to reuse the offline dataset together with newly collected online transitions. One approach initializes the replay buffer with the entire offline dataset, after which new online experiences are appended as the agent interacts with the environment. In this case, the ratio of offline to online data is determined by the dataset size and gradually shifts toward online data as training progresses. An alternative approach maintains two separate replay buffers: one fixed buffer containing the offline dataset and another buffer for online experiences. During training, each batch is sampled from both buffers according to a to a specified offline data ratio $\alpha$. For instance, RLPD (Ball et al., 2023) uses $\alpha=0.5$, corresponding to a symmetric 50% offline and 50% online sampling ratio.
@@ 4.2 Stability with respect to the pretrained policy $\pi_{0}$
Online data warm-up.
Before applying gradient updates, the agent first collects a larger amount of online data ($K$ steps). This strategy reduces the mismatch between the pretraining distribution and the online data distribution, lowering the chance that early updates overwrite prior knowledge. It was introduced in WSRL (Zhou et al., 2024).
Offline RL regularization.
Fine-tuning can also reuse the same offline RL algorithm that produced the pretrained policy. This carries over the conservative regularization used in the offline phase, helping to control distribution shift and overestimation during online updates. Such regularization during fine-tuning has been widely adopted in prior work (Nair et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2023; Nakamoto et al., 2023).
@@ 4.3 Plasticity: Parameter Reset
The most direct way to increase plasticity is to periodically reset network parameters while retaining the accumulated training data. Randomly initialized networks tend to exhibit higher plasticity than pretrained ones (Nikishin et al., 2022), although this strategy may come at the cost of severe forgetting and performance degradation. In the context of offline-to-online RL, parameter reset can be interpreted as starting from any pretrained agent $\pi_{0}$ and then resetting its weights to a randomly initialized agent $\pi_{1}$. While this approach severely degrades initial performance (i.e., $J(\pi_{1})-J(\pi_{0})$ is highly negative), it can significantly enhances plasticity. RLPD (Ball et al., 2023) directly trains an online RL agent from random initialization, which can be viewed as pretraining followed by reset before fine-tuning.
@@ 5 Empirical Study
We study a diverse set of benchmark tasks and pretraining dataset compositions, following the experimental protocols of prior work (Nakamoto et al., 2023; Zhou et al., 2024). Specifically, we include MuJoCo locomotion, AntMaze, Kitchen, and Adroit domains from D4RL (Fu et al., 2020), covering a total of 21 dataset-task compositions. All experiments are conducted with 10 random seeds to ensure statistical reliability.
Offline Pretraining Phase.
For the offline pretraining phase, we employ two representative offline RL algorithms, Calibrated Q-Learning (CalQL) (Nakamoto et al., 2023) and ReBRAC (Tarasov et al., 2023), as well as an additional deterministic behavior cloning (BC) (Schaal, 1996) pretraining, a commonly used imitation learning algorithm that serves as a simple alternative to offline RL. Notably, BC typically achieves performance comparable to that of the dataset itself. Combining the 21 dataset-task compositions with these 3 pretraining algorithms yields 63 experimental settings in total. Each setting is defined by a specific combination of pretraining algorithm, dataset, and task. In this work, since we focus on cases where the offline dataset and the task are from the same MDP, the pretraining algorithm and dataset are sufficient to uniquely specify a setting.
@@
Online Fine-Tuning Phase.
To ensure consistency between offline pretraining phase and online fine-tuning phase, we fine-tune each agent using the corresponding base algorithm. Specifically, we fine-tune CalQL-pretrained agents using SAC (Haarnoja et al., 2018) and ReBRAC-pretrained agents using TD3 (Fujimoto et al., 2018). For the deterministic BC pretraining, we use TD3 for fine-tuning to match its deterministic actor structure.
@@
  • Baseline: Fine-tunes the pretrained policy using an online RL algorithm with only online data.
  • $\bm{\pi_{0}}$-centric methods: Two variants of such methods are evaluated: the baseline with (i) online data warm-up (K=5, [^0] K=5{,}000 steps) and (ii) offline RL regularization using the pretraining coefficient.
@@
  • Mixed $\bm{\pi_{0}+\mathcal{D}}$ methods: the baseline combined with offline data replay and offline RL regularization, a combination widely adopted in prior work (Nair et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2023; Nakamoto et al., 2023).
In each setting, we focus and compare the strongest $\pi_{0}$-centric and $\mathcal{D}$-centric methods to better approximate the ideal performance achievable by each stability source, while minimizing confounding from implementation details and hyperparameter tuning. The following subsections present and analyze the fine-tuning results across the three regimes.
@@ 5.1 Superior regime: $J(\pi_{0})>J(\pi_{\mathcal{D}})$
In the Superior regime, the pretrained policy $\pi_{0}$ achieves substantially higher performance than the behavior policy underlying the offline dataset, i.e., $J(\pi_{0})>J(\pi_{\mathcal{D}})$. In such cases, the offline dataset offers limited additional value, and the primary concern becomes preserving stability relative to the pretrained policy during online fine-tuning.
@@
Representative results are shown in Figure 3, with the complete set of results for all settings in this regime provided in the appendix (Figure 6). We perform $t$-tests between the strongest $\pi_{0}$-centric and $\mathcal{D}$-centric methods in each setting within this regime. The results indicate $\pi_{0}$-centric methods outperform $\mathcal{D}$-centric methods in 24 out of 32 settings (75%), while the remaining settings mostly show no statistically significant difference. Only two settings exhibit the opposite outcome, and in both cases the difference in average performance is negligible. One such case is shown in the rightmost plot of the second row in Figure 3. These statistics correspond to the Superior column of the confusion matrix in Table 1.
This aggregate outcome strongly supports our principle that, in the Superior regime, $\pi_{0}$-centric methods tend to be more effective than $\mathcal{D}$-centric methods. In other words: when the pretrained policy already outperforms the dataset (the superior case), methods that stick close to the pretrained policy work better than those that keep leaning on the offline dataset. While the prediction accuracy is not perfect, such discrepancies are anticipated given the influence of hyperparameters and implementation details. Importantly, the overall observed patterns remain consistent with our principle.
Beyond aggregate comparisons, the analysis of specific design choices highlights key trade-offs between stability and plasticity. Within the two variants of $\pi_{0}$-centric methods, online data warm-up achieves better performance in 27 out of 32 settings, reflecting its ability to preserve the pretrained policy’s knowledge while maintaining sufficient plasticity for fine-tuning. In contrast, offline RL regularization provides stronger stability, leading to substantially less performance degradation during the early fine-tuning phase. However, this strong stability comes at the cost of reduced plasticity, which limits long-term improvements during fine-tuning, making it better than online data warm-up in only 5 out of 32 settings, where the pretrained policy is already close to the optimal policy. The same applies to the combination of offline RL regularization with offline data replay, which exhibits the strongest stability among all methods considered. These contrasts highlight the importance of considering each setting and identifying the the method that best balances the underlying stability–plasticity trade-off.
@@
5.2 Inferior Regime: $J(\pi_{0})<J(\pi_{\mathcal{D}})$
In the Inferior regime, the pretrained policy $\pi_{0}$ performs substantially worse than the behavior policy underlying the offline dataset, i.e., $J(\pi_{0})<J(\pi_{\mathcal{D}})$. In this regime, the pretrained policy contributes substantially less useful knowledge than the offline dataset, making it crucial to retain and leverage the offline data during online fine-tuning.
Representative results are shown in Figure 4, with the complete set of results for all settings in this regime provided in the appendix (Figure 7). According to the results, $\mathcal{D}$-centric methods outperform $\pi_{0}$-centric methods in 19 out of 23 settings (83%), while the remaining settings mostly show no statistically significant difference. Only one setting exhibits the opposite outcome, which is illustrated in the rightmost plot of the second row in Figure 4. These statistics correspond to the Inferior column of the confusion matrix in Table 1. Taken together, these aggregate results show that in the Inferior regime, $\mathcal{D}$-centric methods tend to be more effective than $\pi_{0}$-centric methods, consistent with the prediction of our framework.
@@ 5.3 Comparable Regime: $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$
In the Comparable regime, the pretrained policy and the dataset behavior policy achieve similar performance, i.e., $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$. Representative results are shown in Figure 5, with the complete set of results provided in the appendix (Figure 8).
@@ 6 Conclusion
Our regime taxonomy provides an efficient lens for understanding offline-to-online RL by compressing complex phenomena into a small number of discrete categories. Such taxonomies have been highly successful in both the natural sciences and machine learning, for example in imitation learning, where regimes have been proposed based on density ratios (Spencer et al., 2021) or dataset size (Belkhale et al., 2023). At the same time, we recognize that discretizing behavior into three regimes is a simplification. Real systems often lie along a continuum, and other dataset characteristics—such as coverage—play an important role but remain difficult to capture consistently. Extending the framework to incorporate such dimensions is an important direction for future work.
@@ Appendix A Related Work
Offline-to-online RL seeks to combine the strengths of offline pretraining with the adaptability of online fine-tuning. Early approaches focused on extending offline RL regularization methods into the online regime, constraining fine-tuning updates to remain close to the pretrained policy. For example, Advantage Weighted Actor-Critic (AWAC) (Nair et al., 2020) and Implicit Q-Learning (IQL) (Kostrikov et al., 2021) applied offline regularization techniques directly to online fine-tuning. Building on the similar idea, PROTO (Li et al., 2023) introduced KL regularization to explicitly constrain the online policy to the pretrained one. Another line of work proposes new replay strategies for incorporating offline data more effectively. Lee et al. (2022) propose balanced replay to mitigate distribution shift and bootstrap error when transitioning from offline to online learning, Liu et al. (2024) employs a diffusion model to select or generate samples. Alternative strategies separate the roles of exploration and exploitation during fine-tuning. Jump-Start RL (JSRL) (Uchendu et al., 2023), maintains a fixed guided policy from pretraining alongside an exploration policy that is updated online, progressively transferring control from the pretrained policy to the learned one. More recent directions include expanding the action space via policy set expansion (PEX (Zhang et al., 2023)) and Bayesian methods for uncertainty-aware exploration (BOORL (Hu et al., 2024)). Some approaches skip offline pretraining but still make use of offline data. For example, Song et al. (2022) and Ball et al. (2023) start training directly with online RL while incorporating offline datasets, providing another way to combine offline data with online interaction.
@@
Plasticity and stability have long been recognized as central, often competing, objectives in learning systems. In neuroscience, this tension is formalized as the plasticity–stability dilemma (Mermillod et al., 2013), highlighting the challenge of integrating new knowledge without overwriting previously acquired competencies. In machine learning, similar dynamics manifest when agents must adapt to new data while preserving useful prior knowledge. Early work on continual and lifelong learning addressed this challenge via regularization techniques (Kirkpatrick et al., 2017), replay buffers (Rolnick et al., 2019) and modular architectures (Rusu et al., 2016). More recently, researchers have observed that insufficient plasticity can also hinder online deep RL, motivating methods designed to enhance plasticity of the neural network during training (Nikishin et al., 2022; Sokar et al., 2023; Dohare et al., 2024). In our work, we extend this perspective by framing offline-to-online RL explicitly as a stability–plasticity problem, where the central challenge is to balance the preservation of prior knowledge from pretraining with the adaptability needed to learn from new online experiences.

### v2

@@ 4.1 Stability Relative to the Offline Dataset $\mathcal{D}$
A common strategy for incorporating offline data during fine-tuning is to reuse the offline dataset together with newly collected online transitions. One approach initializes the replay buffer with the entire offline dataset, after which new online experiences are appended as the agent interacts with the environment. In this case, the ratio of offline to online data is determined by the dataset size and gradually shifts toward online data as training progresses. An alternative approach maintains two separate replay buffers: one fixed buffer containing the offline dataset and another buffer for online experiences. During training, each batch is sampled from both buffers according to a specified offline data ratio $\alpha$. For instance, RLPD (ball2023efficient) uses $\alpha=0.5$, corresponding to a symmetric 50% offline and 50% online sampling ratio.
@@ 4.2 Stability with Respect to the Pretrained Policy $\pi_{0}$
Online data warmup. Before applying gradient updates, the agent $\pi_{0}$ first collects a larger amount of online data ($K$ steps). This strategy reduces the mismatch between the pretraining distribution and the online data distribution, lowering the chance that early updates overwrite prior knowledge. It was introduced in WSRL (zhou2024efficient).
Offline RL regularization. Fine-tuning can also reuse the same offline RL algorithm that produced the pretrained policy, thereby inheriting its conservative regularization. This regularization penalizes state-action pairs outside the online data. Since the online data is collected by the sequence of policies from $\pi_{0}$ to $\pi_{N}$, the regularization implicitly anchors learning around the region visited by $\pi_{0}$, even if $\pi_{0}$ is not stored during fine-tuning. When offline data $\mathcal{D}$ is also used with regularization, we consider it as promoting stability towards both $\pi_{0}$ and $\mathcal{D}$; since $\pi_{0}$ is derived from $\mathcal{D}$, this setting tends to have the strongest stability. Such regularization is widely adopted in prior work (nair2020awac; kostrikov2021offline; tarasov2023revisiting; nakamoto2023cal).
@@ 4.3 Plasticity: Parameter Reset
A direct way to increase plasticity is to periodically reset network parameters while retaining the accumulated training data. Randomly initialized networks tend to exhibit higher plasticity than pretrained ones (nikishin2022primacy), although this strategy may come at the cost of severe forgetting and performance degradation. In the context of offline-to-online RL, parameter reset can be interpreted as starting from any pretrained agent $\pi_{0}$ and then resetting its weights to a randomly initialized agent $\pi_{1}$. While this approach severely degrades initial performance (i.e., $J(\pi_{1})-J(\pi_{0})$ is highly negative), it can significantly enhances plasticity. RLPD (ball2023efficient) directly trains an online RL agent from random initialization, which can be viewed as pretraining followed by fully parameter reset before fine-tuning.
@@ 5 Empirical Study
We study a diverse set of benchmark tasks and pretraining dataset compositions, following the experimental protocols of prior work (nakamoto2023cal; zhou2024efficient). Specifically, we include MuJoCo locomotion, AntMaze, Kitchen, and Adroit domains from D4RL (fu2020d4rl), covering a total of 21 dataset-task compositions. All experiments are conducted with 10 random seeds to ensure statistical reliability.
Offline pretraining phase. We employ two representative offline RL algorithms, Calibrated Q-Learning (CalQL) (nakamoto2023cal) and ReBRAC (tarasov2023revisiting), as well as behavior cloning (BC) (schaal1996learning) using a deterministic policy. To pair a behavior-cloned policy with a critic, we pretrain the critic by Fitted Q Evaluation (FQE) (voloshin2019empirical) after BC. Combining the 21 dataset-task compositions with these 3 pretraining algorithms yields 63 experimental settings in total. Each setting is defined by a specific combination of pretraining algorithm, dataset, and task. In this work, since we focus on cases where the offline dataset and the task are from the same MDP, the pretraining algorithm and dataset are sufficient to uniquely specify a setting.
@@
Online fine-tuning phase. To ensure consistency between offline pretraining phase and online fine-tuning phase, we fine-tune each agent using the corresponding base algorithm. Specifically, we fine-tune CalQL-pretrained agents using SAC (haarnoja2018soft) and ReBRAC-pretrained agents using TD3 (fujimoto2018addressing). For the deterministic BC pretraining, we use TD3 for fine-tuning to match its deterministic actor structure, and use ReBRAC when applying regularization.
@@
  • Baseline: Fine-tuning the pretrained policy using an online RL algorithm with only online data.
  • $\bm{\pi_{0}}$-centric methods: Two variants of such methods are evaluated: the baseline with (i) online data warmup (K=5, [^0] K=5{,}000 steps) and (ii) offline RL regularization using the pretraining coefficient.
@@
  • Mixed $\bm{\pi_{0}+\mathcal{D}}$ methods: The baseline combined with offline data replay and offline RL regularization, a combination widely adopted in prior work (nair2020awac; kostrikov2021offline; tarasov2023revisiting; nakamoto2023cal).
In each setting, we focus and compare the strongest $\pi_{0}$-centric and $\mathcal{D}$-centric methods to better approximate the ideal performance achievable by each stability source, while minimizing confounding from implementation details and hyperparameter tuning. In the following subsections, we first present the empirical values of stability and plasticity across the three regimes and discuss how they align with intuition. We then present and analyze the fine-tuning results across the three regimes. Finally, we examine how the Q-function evolves during fine-tuning in the different regimes.
5.1 Empirical Values of Stability and Plasticity
Table 1: Empirical values of stability, plasticity, and improvement (stability + plasticity) for different fine-tuning methods in the Superior and Inferior regimes (500k environment steps).
Fine-tuning method Stability $\uparrow$ Plasticity $\uparrow$ Improvement $\uparrow$
Superior Regime (500k environment steps)
baseline $0.399\pm 0.363$ $0.832\pm 0.275$ $0.433\pm 0.276$
+ warmup ($\pi_{0}$-centric) $0.394\pm 0.371$ $0.865\pm 0.251$ $\mathbf{0.471\pm 0.258}$
+ offline RL ($\pi_{0}$-centric) $0.199\pm 0.268$ $0.519\pm 0.291$ $0.319\pm 0.256$
+ offline data ($\mathcal{D}$-centric) $0.376\pm 0.380$ $0.796\pm 0.282$ $0.421\pm 0.251$
+ offline data + reset ($\mathcal{D}$-centric) $0.615\pm 0.338$ $\mathbf{1.011\pm 0.183}$ $0.396\pm 0.271$
+ offline RL + offline data (mixed $\pi_{0}+\mathcal{D}$) $\mathbf{-0.124\pm 0.213}$ $0.393\pm 0.288$ $0.270\pm 0.233$
Inferior Regime (500k environment steps)
baseline $0.692\pm 0.307$ $0.475\pm 0.378$ $0.217\pm 0.480$
+ warmup ($\pi_{0}$-centric) $0.677\pm 0.306$ $0.495\pm 0.389$ $0.182\pm 0.490$
+ offline RL ($\pi_{0}$-centric) $0.650\pm 0.322$ $0.262\pm 0.283$ $0.388\pm 0.489$
+ offline data ($\mathcal{D}$-centric) $0.674\pm 0.308$ $696\pm 0.340 0.0$$023\pm 0.451$
+ offline data + reset ($\mathcal{D}$-centric) $0.769\pm 0.275$ $\mathbf{0.870\pm 0.299}$ $\mathbf{0.101\pm 0.205}$
+ offline RL + offline data (mixed $\pi_{0}+\mathcal{D}$) $\mathbf{-0.614\pm 0.354}$ $0.440\pm 0.297$ $0.174\pm 0.441$
@@ 5.1 Empirical Values of Stability and Plasticity
To ground the concepts in knowledge decomposition (Eq. 6) in practice, we report the empirical values of stability, plasticity, and the improvement for different fine-tuning methods in Table 1 (Superior and Inferior regimes), Table 7 (Comparable regime), and Table 8 (all regimes).
We observe that the empirical values of stability and plasticity align well with intuition:
  • Stability and Plasticity: The “offline RL + offline data” method attains the highest average stability across all regimes, as it improves stability relative to both $\pi_{0}$ and $\mathcal{D}$. It also achieves the highest plasticity in the Comparable regime. In contrast, “offline data + reset” yields the highest average plasticity in the Superior and Inferior regimes, but this comes at the cost of driving stability to its lowest possible level due to the full parameter reset. We also provide additional analysis in B.2.
  • Improvement: The highest average improvement is regime-dependent. In the Superior regime, it is achieved by “warmup”, which is a $\pi_{0}$-centric method; in the Inferior regime, it is achieved by “offline data + reset”, which is a $\mathcal{D}$-centric method. In the Comparable regime, the best-performing method is “offline RL + offline data”, with a small advantage. These patterns are consistent with intuition: in the Superior regime, $\pi_{0}$ provides more valuable knowledge to leverage, whereas in the Inferior regime, the offline dataset $\mathcal{D}$ contains the more valuable knowledge.
  • Across versus within regimes: Aggregated results (Table 8) show modest differences across methods, but regime-specific tables reveal clear, method-dependent patterns, highlighting the practical utility of the stability–plasticity principle and the three-regime framework.
@@ 5.2 Superior Regime: $J(\pi_{0})>J(\pi_{\mathcal{D}})$
In this regime, the pretrained policy $\pi_{0}$ achieves substantially higher performance than the behavior policy, i.e., $J(\pi_{0})>J(\pi_{\mathcal{D}})$. In such cases, the offline dataset offers limited additional value, and the primary concern becomes preserving stability relative to $\pi_{0}$ during online fine-tuning.
@@
Representative results are shown in Figure 3, with the complete results in this regime provided in the appendix (Figure 7). We perform $t$-tests between the strongest $\pi_{0}$-centric and $\mathcal{D}$-centric methods in each setting. The results indicate $\pi_{0}$-centric methods outperform $\mathcal{D}$-centric methods in 24 out of 32 settings (75%), while the remaining settings mostly show no statistically significant difference. Only two settings exhibit the opposite outcome, and in both cases the difference in average performance is negligible. One such case is shown in the bottom-right plot of Figure 3. These statistics correspond to the Superior column of the confusion matrix in Table 2.
These aggregate outcomes strongly support our principle that, in the Superior regime, $\pi_{0}$-centric methods tend to be more effective than $\mathcal{D}$-centric methods. In other words, when the pretrained policy $\pi_{0}$ already outperforms the dataset (the superior case), methods that stick close to $\pi_{0}$ work better than those that keep leaning on the offline dataset. While the prediction accuracy is not perfect, such discrepancies are anticipated given the influence of hyperparameters and implementation details. Importantly, the overall observed patterns remain consistent with our principle.
Beyond aggregate comparisons, the analysis of specific design choices highlights key trade-offs between stability and plasticity. Within the two variants of $\pi_{0}$-centric methods, online data warmup achieves better performance in 27 out of 32 settings, reflecting its ability to preserve the pretrained policy’s knowledge while maintaining sufficient plasticity for fine-tuning. In contrast, offline RL regularization provides stronger stability, leading to substantially less performance degradation during the early fine-tuning phase. However, this strong stability comes at the cost of reduced plasticity, which limits long-term improvements during fine-tuning, making it better than online data warmup in only 5 out of 32 settings, where the pretrained policy is already close to the optimal policy. The same applies to the combination of offline RL regularization with offline data replay, which exhibits the strongest stability among all methods considered. These contrasts highlight the importance of considering each setting and identifying the the method that best balances the underlying stability–plasticity trade-off.
@@
5.3 Inferior Regime: $J(\pi_{0})<J(\pi_{\mathcal{D}})$
In this regime, the pretrained policy $\pi_{0}$ performs much worse than the behavior policy underlying the offline dataset, i.e., $J(\pi_{0})<J(\pi_{\mathcal{D}})$. Thus, $\pi_{0}$ contributes substantially less useful knowledge than the offline dataset, making it crucial to retain and leverage the offline data during online fine-tuning.
@@ 5.3 Inferior Regime: $J(\pi_{0})<J(\pi_{\mathcal{D}})$
Representative results are shown in Figure 4, with the complete results in this regime provided in the appendix (Figure 8). $\mathcal{D}$-centric methods outperform $\pi_{0}$-centric methods in 19 out of 23 settings (83%), while the remaining settings mostly show no statistically significant difference. Only one setting exhibits the opposite outcome, which is illustrated in the rightmost plot of the second row in Figure 4. These statistics correspond to the Inferior column of the confusion matrix in Table 2. Taken together, these aggregate results show that in the Inferior regime, $\mathcal{D}$-centric methods tend to be more effective than $\pi_{0}$-centric methods, consistent with the prediction of our framework.
@@ 5.4 Comparable Regime: $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$
In this regime, the pretrained policy and the behavior policy achieve similar performance, i.e., $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$. Representative results are shown in Figure 5, with the complete results provided in the appendix (Figure 9).
@@
5.5 Analysis of Q-Function During Fine-Tuning
Lastly, because critic learning is central to value-based methods, we examine how the Q-function evolves during fine-tuning in the Superior and Inferior regimes. Figure 6 illustrates a representative pattern using CalQL pretraining, with one Superior regime (kitchen-partial-v0) and one Inferior regime (kitchen-complete-v0). In each regime, we compare a $\pi_{0}$-centric method (warmup) with a $\mathcal{D}$-centric method (offline data).
In both regimes, the two methods exhibit similarly low TD loss on the online data distribution. As expected, fine-tuning with the offline dataset results in small TD loss on the offline dataset. However, warmup (without offline dataset) produces much larger TD loss on the offline dataset in the Inferior regime than in the Superior regime, especially in the first 100k steps, and the corresponding Q-values diverge more severely. This further illustrates why the offline dataset is crucial in the Inferior regime, whereas it is not necessary in the Superior regime.
Figure 6: Comparison of Q-function behavior in different regimes. Fine-tuning without offline data results in significantly higher TD loss on the offline dataset in the Inferior regime compared to the Superior regime, and also leads to divergence of the Q-values on the offline dataset.
@@ 6 Conclusion
Our regime taxonomy provides an efficient lens for understanding offline-to-online RL by compressing complex phenomena into a small number of discrete categories. Such taxonomies have been highly successful in both the natural sciences and machine learning, for example in imitation learning, where regimes have been proposed based on density ratios (spencer2021feedback) or dataset size (belkhale2023data). At the same time, we recognize that discretizing behavior into three regimes is a simplification, particularly in situations where return does not adequately capture the usefulness of pretrained policies or offline datasets, for example in long-horizon sparse-reward settings. Real systems often lie along a continuum, and other dataset characteristics, such as coverage, play an important role but remain difficult to capture consistently. Extending the framework to incorporate such dimensions is an important direction for future work.
@@ Appendix A Related Work
Offline-to-online RL seeks to combine the strengths of offline pretraining with the adaptability of online fine-tuning. Early approaches focused on extending offline RL regularization methods into the online regime, constraining fine-tuning updates to remain close to the pretrained policy. For example, Advantage Weighted Actor-Critic (AWAC) (nair2020awac) and Implicit Q-Learning (IQL) (kostrikov2021offline) applied offline regularization techniques directly to online fine-tuning. Building on the similar idea, PROTO (li2023proto) introduced KL regularization to explicitly constrain the online policy to the pretrained one. Another line of work proposes new replay strategies for incorporating offline data more effectively. lee2022offline propose balanced replay to mitigate distribution shift and bootstrap error when transitioning from offline to online learning, liu2024energy employs a diffusion model to select or generate samples. Alternative strategies separate the roles of exploration and exploitation during fine-tuning. Jump-Start RL (JSRL) (uchendu2023jump), maintains a fixed guided policy from pretraining alongside an exploration policy that is updated online, progressively transferring control from the pretrained policy to the learned one. More recent directions include expanding the action space via policy set expansion (PEX (PEX)) and Bayesian methods for uncertainty-aware exploration (BOORL (hu2024bayesian)). Some approaches skip offline pretraining but still make use of offline data. For example, song2022hybrid and ball2023efficient start training directly with online RL while incorporating offline datasets, providing another way to combine offline data with online interaction.
@@
Plasticity and stability have long been recognized as central, often competing, objectives in learning systems. In neuroscience, this tension is formalized as the stability–plasticity dilemma (mermillod2013stability), highlighting the challenge of integrating new knowledge without overwriting previously acquired competencies. In machine learning, similar dynamics manifest when agents must adapt to new data while preserving useful prior knowledge. Early work on continual and lifelong learning addressed this challenge via regularization techniques (kirkpatrick2017overcoming), replay buffers (rolnick2019experience) and modular architectures (rusu2016progressive). More recently, researchers have observed that insufficient plasticity can also hinder online deep RL, motivating methods designed to enhance plasticity of the neural network during training (nikishin2022primacy; sokar2023dormant; dohare2024loss). In our work, we extend this perspective by framing offline-to-online RL explicitly as a stability–plasticity problem, where the central challenge is to balance the preservation of prior knowledge from pretraining with the adaptability needed to learn from new online experiences.